OpenStudy (anonymous):

What are the x-intercept(s) of the graph of the equation y − x^2 = −4x − 5?

OpenStudy (campbell_st):

rewrite the equation \[y = x^2 -4x - 5\] to find the x-intercepts let y = 0 then you have \[0 = x^2 - 4x - 5\] this quadratic can be factorised find the factors of -5 that add to -4 the bigger factor is negative and the smaller is positive

OpenStudy (anonymous):

(1, 0) and (−5, 0)..

OpenStudy (campbell_st):

just a bit quick... -5 and 1 are the essential numbers so you factorise and get (x -5)(x+1) = 0 now the intercepts will be the same numbers with opposite signs

OpenStudy (campbell_st):

the reason is that the equation must equal zero I'll substitute x = 5 into the factorised equation y = ( 5 -5)(5 +1) this gives 0 * 6 = 0 same when you subdtitute x = -1 y = (-1 -5)(-1+1) -6 * 0 y = 0

OpenStudy (anonymous):

(−1, 0) and (5, 0)

OpenStudy (campbell_st):

thats correct... there are a few tips about factorising and factors in a quadratic ax^2 + bx + c if b and c are positive both factors are positive b negative and c positive both factors are negative if c is negative 1 factor is negative... b positive the larger factor is positive b negative the larger factor is negative hope it all helps